Solve for x as a function of t. Answers must be exact, not decimal.
ln(x+4)+ln(x-4)=9t

Answer by stanbon(48535)

ln(x+4)+ln(x-4)=9t
=ln[(x+4)(x-4)] = 9t
= ln[(x^2-16)] = 9t
e^(9t) = x^2-16
x^2 = 16+e^9t
x = [sqrt(16+e^9t)] or x = [-sqrt(16+e^9t]
==============
Cheers,
stan H.